# Geological and Well prior assisted full waveform inversion using conditional diffusion models

### v1

@@ 1 Introduction
Our industry has long strived to integrate the seismic value chain, empowering seismic interpreters to play a more significant role in guiding the velocity model-building process. Their expertise in defining geological features, such as faults, horizons, and lithology, is fundamental for accurate model construction. As a result, we have increasingly integrated advanced technology, which can utilize that expertise, into our workflows to enhance the velocity model-building process from seismic data. We have developed specialized software that enables interpreters to interactively analyze seismic volumes, mark geological features, and validate the accuracy of inverted models. However, these approaches are typically post-inversion and aimed at providing guidance for corrections. They are often not incorporated within the inversion process itself, as is the case with Full Waveform Inversion (FWI). Although FWI can reconstruct the subsurface velocity model by matching the simulated seismograms with the observed recordings, its performance on realistic data is not as good as expected often due to imperfect observations, resulting in band limited, noisy, and sparse data. Incorporating the expertise of geologists and other prior information into FWI is crucial for guiding the inversion toward realistic, high-resolution models.
@@
From the perspective of ill-posed inverse problems, the prior information can be incorporated via regularization. In addition to standard (fixed) regularization methods, such as Tikhonov (Aster et al., 2011) and total variation (TV) regularization (Alkhalifah et al., 2018; Xiang and Zhang, 2016), which may yield smooth or piece-wise smooth inverted results, respectively, using deep learning to incorporate prior information, like our expectations on geological structures, migrated images, or well information, to guide the inversion has recently gained attention. Zhang et al. (2018) proposed to utilize the facies-based constraints inverted from Bayesian theory to improve multiparameter elastic FWI. Singh et al. (2018) also proposed to use facies distribution and the available well logs to derive the prior model as a constraint in the inversion workflow. Li et al. (2021) and Zhang and Alkhalifah (2022) proposed the use of a deep neural network to learn the statistical relationship between the inverted model and the facies interpreted from well logs for inversion. In addition to these explicit ways, an effective and implicit choice is to use generative models that can build the data distribution, e.g., generative adversarial network (GAN), trained on a velocity distribution based on our expectations to regularize FWI Mosser et al. (2020). However, we know that it is difficult to train GANs and may face the challenge of model collapse. Diffusion models, which show state-of-art performance compared to other generative models in generative quality and diversity on image generation tasks Dhariwal and Nichol (2021), have gained a lot of attention lately. In addition, the velocity model size nature of the latent space in diffusion models allowed for the seamless integration of pre-trained (with prior information) diffusion models into FWI by means of plug-and-play Wang et al. (2023). As demonstrated by Wang et al. (2023), for FWI, a trained diffusion model can inject prior information of the trained velocity distribution into the inversion results by updating the velocity model in a regularized form after each iteration, yielding the improvement in the accuracy of FWI in case of noisy and sparse observations, referred to as diffusionFWI. Although it showed promising performance in field applications, its prior information is guided by the velocity distribution and does not offer a mechanism to inject specific observed prior or high-level expertise from geologists, such as geological preferences and well information, during the inversion. As mentioned earlier, this information is also crucial to enhance the inversion accuracy and is worth exploring. Revisiting the diffusionFWI, it can be regarded as a conditional sampling process with the recorded surface data and the governing equation of wave propagation as conditions. Hence, other additional prior information, such as well logs or geological structure information, can also be considered as conditions for the diffusionFWI. Wang et al. (2024) demonstrated that using a conditional diffusion model can generate velocity models that satisfy the velocity distribution while also matching the given conditions, including class labels, well logs, and structural information.
@@
To this end, we propose conditional diffusion regularized FWI using conditional diffusion models to inject, for example, geological or well information into the inversion process. The proposed method makes the inverted results not only satisfy the observed data and velocity distribution, but also match the additional information like well logs and geological descriptions. Specifically, we consider the geological label and the well log and its position, which are commonly available in real applications, in the diffusion models’ training as conditions. Then, during the iterative inversion process, these prior information can be implicitly embedded into FWI by applying this conditional diffusion model. To the best of our knowledge, this is the first framework that allows the integration of a geologist’s expertise, such as geological labels (e.g., flatness, curvature, and faults), along with regularization from measurements like well log data—something that was not achievable with conventional regularization techniques or deep learning-based methods. To summarize, our main contributions are as follows:
@@
  • We evaluate the proposed method on synthetic data corresponding to the OpenFWI velocity models and field marine data and demonstrate the effectiveness of the proposed method.
@@ 2.2 Training and inference of diffusion models
where $\lambda$ is the conditioning scale to adjust the strength of the conditioning. A higher value for $\lambda$ will decrease the diversity of the samples but increase the consistency between the higher quality (high-fidelity) sample and the given condition (Ho and Salimans, 2022). To make the diffusion model more flexible and avoid separate training of two neural networks, in training, we randomly add embedding vectors of the condition as well as the vectors of a null condition denoting no conditioning, yielding flexible generation abilities when sampling. With this trick in the training stage, the unconditioned $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{i})$ can be realized by setting $\boldsymbol{c}$ as a vector of null condition in $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{i},\boldsymbol{c})$.
Figure 1: Conditional generative diffusion models. The backbone network is U-net (Ronneberger et al., 2015).$\boldsymbol{x}_{0}$ represents samples from the training models distribution, which are transformed using the Diffusion process to $\boldsymbol{x}_{T-1}$, which are samples corresponding to a normal distribution, and $\boldsymbol{T}$ is the number of time steps. $\mathbb{MLP}$ stands for Multi-Layer Perceptron.
@@ 3.1 The training of Conditional diffusion models on OpenFWI dataset
Here, we train the conditional diffusion model on the OpenFWI dataset to showcase the flexibility of the proposed method to control the velocity generation, specifically with class labels and well logs. The dataset comprises eight distinct classes: "FlatVel-A, "FlatVel-B", "CurveVel-A", "CurveVel-B", "FlatFault-A", "FlatFault-B", "CurveFault-A", "CurveFault-B", whose example samples are shown in Figure 3. They exhibit distinct structural features and velocity variations with depth. For example, the class name "FlatVel-" represents layered velocity models, while "CurveVel-" corresponds to velocity models with folding layers. For those two types of velocity models, "A" represents the gradual increase of velocity with depth, while "B" represents random velocity variation with depth. The "FlatFault-" and "CurveFault-" denote the types of velocity models that include faults with horizontal or curved interfaces, respectively. For "FlatFault-" and "CurveFault-", "A" indicates the case with fewer discontinuities, while "B" indicates the case with more discontinuities and more pronounced velocity changes. The original size of the velocity model is 70$\times$70, and we crop them to a model size of 64$\times$64 for ease of training. To handle the distribution of this dataset, we use a standard U-net (Oktay et al., 2018) with attention blocks, and the feature maps resolution of the U-net is given by $[64\times 64,32\times 32,16\times 16,8\times 8]$. We train the diffusion model using an Adam optimizer with a learning rate of 1e-4. The batch size for the training is 1024, and the maximum training iteration is set to 200000. The velocity model class label and the well log (a randomly chosen vertical profile from the velocity model) are passed through a multi-layer perceptron (MLP) and then fused directly into the U-net layers. For the well, we incorporate a positional encoding so that the network learns the position of the well.
@@ 3.2 Synthetic inversion test
We compute synthetic observed data by numerically solving the acoustic wave equation with a time sampling interval of 1 ms using a 15 Hz Ricker wavelet with a total recording length of 1.5 s. We uniformly place 16 receivers and 4 shots on the surface. Considering the relatively high-frequency and limited acquisition, applying conventional FWI admits poor results (Figure 5a). Meanwhile, including the diffusion model introduces high-resolution components, as shown in Figure 5b. As the class label "CurveFaultA" is associated with high-velocity values deep, as well as sinusoidal-shaped curves and fault structures, the velocity values deep are further improved (Figure 5c) courtesy of including the geological class. Alternatively, using the well-log condition, as shown in Figure 5d, helps improve the velocity values at depth even more. Using both the geological class label and well condition together (Figure 5e), the small fault structure is more visible, which can not be achieved by the conventional FWI or regular diffusion FWI. Figure 6 shows the profile comparison at the well location at 0.25 km. We can see that the well condition reasonably constrains the model.
@@
Figure 4: a)The true velocity model, and b) the initial velocity model. The conditional generation results starting from the initial model using c) the geological class label, d) a well located at 0.25 km, and e) the geological label and the well information together.
@@
Figure 5: Inversion results from a) conventional FWI, b) diffusion FWI, and conditional diffusion FWI with c) the geological class label, d) a well located at 0.25 km, and e) the geological class label and the well information together.
Figure 6: A profile comparison at the location of the well at 0.25 km for the inversion results in Figure 5.
@@
In a more challenging scenario, we applied conditional diffusion FWI with a limited and closely-spaced shot array: four shots are placed at 0 km, 0.02 km, 0.04 km, and 0.06 km, which correspond to the far left side of the model, which should cause poor illumination of the right side of the model. The inversion results are presented in Figure 7. As a result, the conventional FWI fails to recover the subsurface velocity, especially on the right side. The same holds for diffusion FWI, but with generally higher resolution. Incorporating the geological class label, as shown in Figure 7c, improved the subsurface structures. On the other hand, adding well-log information (Figure 7d) enhanced the velocity estimation at depth. The well information helps guide the velocity distribution with depth, especially near the well location. Meanwhile, the geological class information has a more global effect on the prediction, especially on the structure (less on the values). Using the geological class and well condition simultaneously, we obtain the inversion result in Figure 7e, which is the closest to the true velocity among the five results.
@@
Figure 7: Inversion results with a limited and closely-spaced shot array from a) conventional FWI, b) diffusion FWI, and conditional diffusion FWI with c) the geological class label, d) a well located at 0.25 km, and e) the geological class label and the well information together.
@@
From the above tests, we observe clear improvements with geological and well log priors-assisted FWI. Unlike the class label, the well log depends on an important factor, which is the location of the well log. To further test and understand the influence of the well location on the diffusion FWI, we select another test (not used in the training) velocity model shown in Figure 8. This velocity model includes a very complex fault structure and strong lateral velocity variations from shallow to deep. It will help us test the well-log information guidance for complex scenarios. The experiment parameters regarding the geometry and source configurations are the same as the test in Figure 5. We conduct conventional FWI, diffusion FWI, and well-conditioned diffusion FWI with the well logs at different locations, whose results are shown in Figure 9. The corresponding profile comparisons for well condition at each position are shown in Figure 10. We noticed that the well-log information can improve the velocity values, especially at depth. However, due to the strong velocity changes in the lateral direction, this improvement is mainly tied to the well location. An improper well-log location will somehow limit its guidance in diffusion-regularized FWI, resulting in relatively poor inversion results. So, this experiment again demonstrates that the best way to condition diffusion FWI in a complex scenario is to include local and global type conditions, specifically well logs and class labels, in our experiment. This makes sense as the more information considered, the better results we should obtain, which is the main motivation of this paper.
@@
Figure 10: Profile comparison for the three well-conditioned diffusion FWI results (a) at 0.25km (Figure 9c), and (b) at 0.40km (Figure 9d) with the true model in Figure 8a, conventional FWI results in Figure 9a, and diffusion FWI result in Figure 9b, respectively.

### v2

@@ 1 Introduction
Our industry has long strived to integrate the seismic value chain, empowering seismic interpreters to play a more significant role in guiding the velocity model-building process. Their expertise in defining geological features, such as faults, horizons, and lithology, is fundamental for accurate model construction. As a result, we have increasingly integrated advanced technology, which can utilize that expertise, into our workflows to enhance the velocity model-building process from seismic data. We have developed specialized software that enables interpreters to interactively analyze seismic volumes, mark geological features, and validate the accuracy of inverted models. However, these approaches are typically post-inversion and aimed at providing guidance for corrections. They are often not incorporated within the inversion process itself, as is the case with Full Waveform Inversion (FWI). Although FWI can reconstruct the subsurface velocity model by matching the simulated seismograms with the observed recordings, its performance on realistic data is not as good as expected, often due to imperfect observations, resulting in band-limited, noisy, and sparse data. Incorporating the expertise of geologists and other prior information into FWI is crucial for guiding the inversion toward realistic, high-resolution models.
@@
From the perspective of ill-posed inverse problems, the prior information can be incorporated via regularization. In addition to standard (fixed) regularization methods, such as Tikhonov (Aster et al., 2011) and total variation (TV) regularization (Alkhalifah et al., 2018; Xiang and Zhang, 2016), which may yield smooth or piece-wise smooth inverted results, respectively, using deep learning to incorporate prior information, like our expectations on geological structures, migrated images, or well information, to guide the inversion has recently gained attention. Zhang et al. (2018) proposed to utilize the facies-based constraints inverted from Bayesian theory to improve multiparameter elastic FWI. Singh et al. (2018) also proposed to use facies distribution and the available well logs to derive the prior model as a constraint in the inversion workflow. Li et al. (2021) and Zhang and Alkhalifah (2022) proposed the use of a deep neural network to learn the statistical relationship between the inverted model and the facies interpreted from well logs for inversion. In addition to these explicit ways, an effective and implicit choice is to use generative models that can build the data distribution, e.g., generative adversarial network (GAN), trained on a velocity distribution based on our expectations to regularize FWI Mosser et al. (2020). However, we know that it is difficult to train GANs and may face the challenge of model collapse. Diffusion models, which show state-of-the-art performance compared to other generative models in generative quality and diversity on image generation tasks Dhariwal and Nichol (2021), have gained a lot of attention lately. In addition, the velocity-model-sized latent space in diffusion models enables seamless integration of pre-trained diffusion models (with prior information) into FWI in a plug-and-play manner (Wang et al., 2023). As demonstrated by Wang et al. (2023), for FWI, a trained diffusion model can inject prior information of the trained velocity distribution into the inversion results by updating the velocity model in a regularized form after each iteration, yielding the improvement in the accuracy of FWI in case of noisy and sparse observations, referred to as diffusionFWI. Although it showed promising performance in field applications, its prior information is guided by the velocity distribution and does not offer a mechanism to inject specific observed prior or high-level expertise from geologists, such as geological preferences and well information, during the inversion. As mentioned earlier, this information is also crucial to enhance the inversion accuracy and is worth exploring. Revisiting the diffusionFWI, it can be regarded as a conditional sampling process with the recorded surface data and the governing equation of wave propagation as conditions. Hence, other additional prior information, such as well logs or geological structure information, can also be considered as conditions for the diffusionFWI. Wang et al. (2024) demonstrated that using a conditional diffusion model can generate velocity models that satisfy the velocity distribution while also matching the given conditions, including class labels, well logs, and structural information.
@@
To this end, we propose conditional diffusion regularized FWI using conditional diffusion models to inject, for example, geological or well information into the inversion process. The proposed method makes the inverted results not only satisfy the observed data and velocity distribution, but also match the additional information like well logs and geological descriptions. Specifically, we consider the geological label and the well log and its position, which are commonly available in real applications, in the diffusion models’ training as conditions. Then, during the iterative inversion process, these prior information can be implicitly embedded into FWI by applying this conditional diffusion model. To the best of our knowledge, this is the first framework that allows the integration of a geologist’s expertise, such as geological labels (e.g., flatness, curvature, and faults), along with regularization from measurements like well log data, something that was not achievable with conventional regularization techniques or deep learning-based methods. To summarize, our main contributions are as follows:
@@
  • We evaluate the proposed method on synthetic data corresponding to the OpenFWI velocity models and field marine data, and demonstrate the effectiveness of the proposed method.
@@ 2.2 Training and inference of diffusion models
where $\lambda$ is the conditioning scale to adjust the strength of the conditioning. In this study, unconditional generation is performed with $\lambda=-1$, while conditional generation uses $\lambda$ values ranging from 1.0 to 20.0. A higher value for $\lambda$ will decrease the diversity of the samples but increase the consistency between the higher quality (high-fidelity) sample and the given condition (Ho and Salimans, 2022). During training, to avoid separate training of two neural networks, we apply the classifier-free guidance strategy by randomly dropping the conditional inputs with a probability of 0.5. Specifically, for each training sample, the conditioning information is either retained or replaced, with equal probability, with a learnable null embedding initialized from random parameters. This enables the model to learn both conditional and unconditional generation behaviors, which can later be combined at inference time for guided sampling. With this trick in the training stage, the unconditioned $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{i})$ can be realized by setting $\boldsymbol{c}$ as a vector of null condition in $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{i},\boldsymbol{c})$.
Figure 1: Conditional generative diffusion models where different conditions, such as geological labels and well-log velocities, are incorporated for velocity prediction. The backbone network is U-Net (Ronneberger et al., 2015).$\boldsymbol{x}_{0}$ represents samples from the training models distribution, which are transformed using the Diffusion process to $\boldsymbol{x}_{T-1}$, which are samples corresponding to a normal distribution, and $\boldsymbol{T}$ is the number of time steps. $\mathbb{MLP}$ stands for Multi-Layer Perceptron.
@@ 3.1 The training of Conditional diffusion models on OpenFWI dataset
Here, we train the conditional diffusion model on the OpenFWI dataset to showcase the flexibility of the proposed method to control the velocity generation, specifically with class labels and well logs. The dataset comprises eight distinct classes: "FlatVel-A, "FlatVel-B", "CurveVel-A", "CurveVel-B", "FlatFault-A", "FlatFault-B", "CurveFault-A", "CurveFault-B", whose example samples are shown in Figure 3. They exhibit distinct structural features and velocity variations with depth. For example, the class name "FlatVel-" represents layered velocity models, while "CurveVel-" corresponds to velocity models with folding layers. For those two types of velocity models, "A" represents the gradual increase of velocity with depth, while "B" represents random velocity variation with depth. The "FlatFault-" and "CurveFault-" denote the types of velocity models that include faults with horizontal or curved interfaces, respectively. For "FlatFault-" and "CurveFault-", "A" indicates the case with fewer discontinuities, while "B" indicates the case with more discontinuities and more pronounced velocity changes. The original size of the velocity model is 70$\times$70, and we crop it to a model size of 64$\times$64 for ease of training. To handle the distribution of this dataset, we use a standard U-Net (Oktay et al., 2018) with attention blocks, and the feature maps resolution of the U-Net is given by $[64\times 64,32\times 32,16\times 16,8\times 8]$. We train the diffusion model using an Adam optimizer with a learning rate of 1e-4. The batch size for the training is 1024, and the maximum training iteration is set to 200000. For stable convergence, we also apply the Exponential Moving Average (EMA) for model parameter updates. The velocity model class label and the well log (a randomly chosen vertical profile from the velocity model) are passed through a multi-layer perceptron (MLP) and then fused directly into the U-Net layers. For the well, we incorporate a positional encoding so that the network learns the position of the well.
@@ 3.2 Synthetic inversion test
We compute synthetic observed data by numerically solving the acoustic wave equation with a time sampling interval of 1 ms using a 15 Hz Ricker wavelet with a total recording length of 1.5 s. We uniformly place 16 receivers and 4 shots on the surface. Considering the relatively high-frequency and limited acquisition, applying conventional FWI admits poor results (Figure 5a). Meanwhile, including the diffusion model introduces high-resolution components, as shown in Figure 5b. As the class label "CurveFaultA" is associated with high-velocity values deep, as well as sinusoidal-shaped curves and fault structures, the velocity values deep are further improved (Figure 5c) courtesy of including the geological class. Alternatively, using the well-log condition, as shown in Figure 5d, helps improve the velocity values at depth even more. Using both the geological class label and well condition together (Figure 5e), the small fault structure is more visible, which can not be achieved by the conventional FWI or regular diffusion FWI. Figure 6 shows the profile comparison at the well location at 0.25 km. We can see that the well condition reasonably constrains the model. Table 1 reports the quantitative performance of different methods in terms of Structural Similarity Index Measure (SSIM), Peak Signal-to-Noise Ratio (PSNR), and relative L2 error, where “Gen” denotes pure generation without inversion, and “CD-FWI” denotes Conditional DiffusionFWI. SSIM evaluates the perceptual similarity between the predicted and reference images, PSNR reflects the reconstruction fidelity in decibels, and the relative L2 error measures the normalized discrepancy in Euclidean norm. It further demonstrates that the more conditions we have, including data, class, and well, the better results we can obtain.
@@
Figure 4: a)The true velocity model, and b) the initial velocity model. The pure conditional generation results starting from the initial model using c) the geological class label, d) a well located at 0.25 km, and e) the geological label and the well information together. Noted that there are no FWI incorporated for those generations. This demonstrates the importance of incorporating the observed data by FWI.
@@
Figure 5: Inversion results from a) conventional FWI, b) diffusion FWI, and conditional diffusion FWI with c) the geological class label, d) a well located at 0.25 km, and e) the geological class label and the well information together. The proposed conditional diffusion FWI outperforms both conventional methods and standard diffusion-based FWI. Moreover, incorporating more conditioning information leads to improved inversion results.
Figure 6: A profile comparison at the location of the well at 0.25 km for the inversion results in Figure 5. Detailed profile comparisons further validate that the proposed conditional diffusion FWI outperforms other approaches. Incorporating well constraints improves the match between the inverted results and well data, while jointly using both well and class conditioning yields the best performance.
Table 1: Quantitative comparison of different approaches using PSNR, SSIM, and Relative L2 error. Higher PSNR and SSIM, lower relative L2 error are better.
[…]
Gen (class) 23.47 0.289 0.0772
Gen (well) 23.44 0.321 0.0835
Gen (class&well) 25.49 0.495 0.0659
Conventional FWI 28.10 0.670 0.0489
Diffusion FWI 36.49 0.900 0.0183
CD-FWI (class) 37.96 0.958 0.0156
CD-FWI (well) 41.22 0.953 0.0108
CD-FWI (class&well) 41.51 0.976 0.0105
@@
In a more challenging scenario, we applied conditional diffusion FWI with a limited and closely-spaced shot array: four shots are placed at 0 km, 0.02 km, 0.04 km, and 0.06 km, which correspond to the far left side of the model, which should cause poor illumination of the right side of the model. The inversion results are presented in Figure 7. As a result, the conventional FWI fails to recover the subsurface velocity, especially on the right side. The same holds for diffusion FWI, but with generally higher resolution. Incorporating the geological class label, as shown in Figure 7c, improved the subsurface structures. On the other hand, adding well-log information (Figure 7d) enhanced the velocity estimation at depth. The well information helps guide the velocity distribution with depth, especially near the well location. Meanwhile, the geological class information has a more global effect on the prediction, especially on the structure (less on the values). Using the geological class and well condition simultaneously, we obtain the inversion result in Figure 7e, which is the closest to the true velocity among the five results. The quantitative statistics in Table 2 further demonstrate that the proposed approach consistently outperforms conventional FWI and diffusion-based FWI methods.
@@
Figure 7: Inversion results with a limited and closely-spaced shot array from a) conventional FWI, b) diffusion FWI, and conditional diffusion FWI with c) the geological class label, d) a well located at 0.25 km, and e) the geological class label and the well information together. Under limited observation conditions, both conventional FWI and standard diffusion-based FWI fail to produce accurate results, whereas the proposed method achieves significantly better performance. The diffusion FWI, conditioned on both well and class information, delivers the most accurate inversion.
Table 2: Quantitative comparison of different approaches with a limited and closely-spaced shot array. Higher PSNR and SSIM, lower relative L2 error are better.
[…]
Conventional FWI 25.31 0.468 0.0736
Diffusion FWI 25.47 0.543 0.0630
CD-FWI (class) 26.83 0.649 0.0586
CD-FWI (well) 27.80 0.623 0.0504
CD-FWI (class&well) 28.15 0.623 0.0485
@@
From the above tests, we observe clear improvements with geological and well log priors-assisted FWI. Unlike the class label, the well log depends on an important factor, which is the location of the well log. To further test and understand the influence of the well location on the diffusion FWI, we select another test (not used in the training) velocity model shown in Figure 8. This velocity model includes a very complex fault structure and strong lateral velocity variations from shallow to deep. It will help us test the well-log information guidance for complex scenarios. The experiment parameters regarding the geometry and source configurations are the same as the test in Figure 5. We conduct conventional FWI, diffusion FWI, and well-conditioned diffusion FWI with the well logs at different locations, whose results are shown in Figure 9, and the corresponding quantitative comparison is shown in Table 3. The corresponding profile comparisons for well condition at each position are shown in Figure 10. We noticed that the well-log information can improve the velocity values, especially at depth. However, due to the strong velocity changes in the lateral direction, this improvement is mainly tied to the well location. An improper well-log location will somehow limit its guidance in diffusion-regularized FWI, resulting in relatively poor inversion results. So, this experiment again demonstrates that the best way to condition diffusion FWI in a complex scenario is to include local and global type conditions, specifically well logs and class labels, in our experiment. This makes sense as the more information considered, the better results we should obtain, which is the main motivation of this paper.
@@
Table 3: Quantitative comparison of different approaches with different conditions (shown in Figure 9). Higher PSNR and SSIM, lower relative L2 error are better.
[…]
Conventional FWI 27.51 0.727 0.0528
Diffusion FWI 29.21 0.786 0.0417
CD-FWI (well at 0.25 km) 29.57 0.788 0.0401
CD-FWI (well at 0.40 km) 29.96 0.810 0.0380
Figure 10: Profile comparison for the three well-conditioned diffusion FWI results (a) at 0.25km (Figure 9c), and (b) at 0.40km (Figure 9d) with the true model in Figure 8a, conventional FWI results in Figure 9a, and diffusion FWI result in Figure 9b, respectively. It highlights the importance of the well location for the accuracy of the well-conditioned diffusion FWI.